OpenStudy (anonymous):

A circle in the standard (x,y) coordinate plane has its center at (-4,3), and an area of 9π square coordinate units. Which of the following is an equation for the circle? A. (x+4)^2-(y-3)^2 =81 B. (x-3)^2+(y+4)^2 =9 C. (x+4)^2+(y-3)^2 =9 D. (x-4)^2+(y-3)^2 =81 E. (x+4)^2+(y-3)^2 =81

OpenStudy (anonymous):

B

OpenStudy (cwrw238):

general form of circle equation is (x -a)^2 + (x-b)^2 = r^2 where (a,b) is the center and r = radius

OpenStudy (anonymous):

Since the equation of a circle is (x-a)^2 + (x-b)^2 = r^2, where (a, b) is the centre, and area = pi r^2, we know that r is 3. So all we have to do is see which of B or C fits with the equation of a circle and we see that it is B.

OpenStudy (cwrw238):

thats better traxter - an explanation of how you got there

jimthompson5910 (jim_thompson5910):

it's not B

jimthompson5910 (jim_thompson5910):

the center is (-4,3), so h = -4 and k = 3

OpenStudy (anonymous):

It is C

OpenStudy (cwrw238):

yes

jimthompson5910 (jim_thompson5910):

The general form is (x-h)^2+(y-k)^2=r^2

OpenStudy (anonymous):

What a shameful mistake

jimthompson5910 (jim_thompson5910):

its ok

OpenStudy (cwrw238):

thats not shameful! - just human

OpenStudy (anonymous):

I mean....i was doing it to test the person asking the question of course! But seriously, this highlights a mistake I see very often in my students. Be very careful not to mix up the x and y co-ordinates. Also realise when you've been online too long and get to bed :P